# Transferable Watermarking to Self-supervised Pre-trained Graph Encoders by Trigger Embeddings

### v1

@@ IV-A3 Watermark Embedding
During encoder pre-training, we pre-train all the encoders using the aforementioned total loss to embed. The factors in the loss function $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are set as $5.0$, $1.0$ and $1.0$ respectively by default. The trigger node set consists 50 ego-graphs (50 randomly sampled center nodes) in default. All the GSSL models are pre-trained using Adam optimizer [20] with initial learning rate $0.01$ for 2000 epochs.
IV-A4 Downstream Tasks.
@@
Node classification. For node classification, we simulate the downstream classifier as a 2-layer multi-layer perceptron (MLP) with 256 neurons in each hidden layer. When training, we simulate the adversary by using 20 labeled nodes per class to train the downstream classifier. In addition to that, 500 nodes are used for validation and 1000 nodes are for testing. Such settings closely follow [3]. As commonly in classification tasks, we use accuracy (ACC) to record the performance of the node classification task.
@@
Community detection. Following the literature of graph representation learning regarding the task of community detection, we first adopt PCA to reduce the dimensionality of node representations. Then, k-means clustering is used to the representation to cluster the node communities. Here, we use normalized mutual information (NMI) for criterion.
TABLE I: The fidelity evaluation (clean model performance |||| watermarked model performance) of the watermarking method
@@
Figure 2: T-SNE visualization of the node embeddings generated by watermarked encoders. Scattered nodes in different colors represent the embeddings of nodes in different categories. Particularly, trigger embeddings are in black. (a) GGD, cora, (b) GGD, citeseer, (c) GraphCL, cora, (d) GraphCL, citeseer
@@ IV-B1 Fidelity
Fidelity means that the impair of performance degradation induced by watermarking should be low. Specifically, for tasks of node classification, edge prediction and community detection, fidelity is reflected by the difference of ACC, AUC and NMI with and without watermarking, respectively. The experimental results of fidelity evaluation are listed in Table 1, where the normal performance of watermarked and non-watermarked models in downstream tasks are listed. It can be seen that for most models in all the downstream tasks, the performance degradation caused by watermarking is negligible, which is reasonable since the utility loss and the watermarking loss are optimized together.
TABLE II: The concentration score ($CS$) of the trigger predictions produced by watermarked models and non-watermarked models ($CS_{\text{wm}}\%|CS_{\text{clean}}\%)$
@@
Figure 3: The watermark performance w.r.t. model pruning with different pruning rates. Here, we record the curves of concentration scores of trigger embeddings (CS) and model normal performance in three downstream tasks (ACC, AUC and NMI). (a) GGD, Cora, (b) GGD, Citeseer
[…]
Factor Settings DGI GGD GraphCL GraphMAE2
$\lambda_{1}$ $\lambda_{2}$ Cora Citeseer Cora Citeseer Cora Citeseer Cora Citeseer
$\times$ $\checkmark$ 19.1 25.2 18.7 23.3 19.8 21.8 23.3 17.9
$\checkmark$ $\times$ 35.2 37.9 37.5 38.8 79.3 82.1 76.9 85.3
$\checkmark$ $\checkmark$ 67.9 62.5 74.9 73.8 74.7 66.5 78.5 72.7
@@ IV-B2 Transferability
Transferability means that watermarked graph encoder can be verified with models integrated with the encoder. In the proposed method, it means that the concentration score of the watermarked model is larger than a promissory threshold $\tau$. In our experiments, we set the $\tau$ to be 0.7, which can well discriminate watermark and clean models in experience. The transferability evaluations are recorded in Table 2, from which we can see that the concentration scores ($CS_{\text{WM}}$) are generally higher than 0.8 for watermarked models, demonstrating the effectiveness of the watermarking method.
@@ IV-B3 Uniqueness
Uniqueness means that the watermark can be extracted from only watermarked models. To show the uniqueness of the watermark, we list the concentration scores of trigger set representations produced by clean models in Table 2. From the table we can observe that, the concentration scores of prediction results produced by watermarked models are higher than that of non-watermarked clean models by a large margin. For watermarked models, the predictions of trigger set generally exhibit high concentration score larger than the threshold $\tau=0.7$. While for non-watermarked models, since such models have not learned to produce similar embeddings for trigger samples, the result distribution is more uniform, which yields low concentration scores.
@@ IV-B4 Robustness
As a primary approach, we mainly focus on testing the robustness against model parameter pruning since it is the most common way to modifying neural networks. Specifically, we record the change of concentration score and normal performance after pruning $\alpha\%$ of parameters with largest altitude. It can seen in Figure 3 that, generally the concentration score of watermarked model degrades simultaneously with the model normal performance of different tasks as the pruning rate increases. The phenomenon indicates that if is hard for the adversary to remove the watermark without degrading the model normal performance.
@@ IV-B5 Visualization
To further explain the effectiveness of the proposed method, we show the t-SNE visualization of the embeddings produced by watermarked encoder in Figure 2. It can be seen that, the normal nodes are well represented by embeddings while the trigger embeddings form a compact cluster (black dots in the figure) distant from normal samples. The t-SNE visualization clearly shows that the design of watermark loss successfully accomplishes our main objective and thus results in effectiveness watermarking for GSSL encoders.
@@ IV-C Ablation Study
To validate the necessity of the modules in the watermarking scheme, we perform ablation study. Since the watermark is embedded by the watermark loss, so we mainly focus on studying the necessity of loss terms, including $L_{\text{in}}$ and $L_{\text{ext}}$, which are scaled by $\lambda_{1}$ and $\lambda_{2}$, respectively. Here, we focus on node classification and calculate the concentration scores of classification results of models with encoders pre-trained with different loss configurations. The results are shown in Table 3. It can be seen that, when we only use $L_{\text{ext}}$, it is difficult for the model to produce similar embeddings for the trigger set, which is quite reasonable. However, it is interesting that, when we only use $L_{\text{in}}$, only the concentration scores of GraphCL and GraphMAE2 are large, while the concentration scores of GraphCL and GraphMAE2 are low. The results indicate that only for some encoders $L_{\text{ext}}$ are necessary. We hypothesize the reason is that, for some models, only using $L_{\text{in}}$ to cluster the trigger embeddings is insufficient and leads to unstable optimization. While seperating the trigger embeddings and normal embeddings may help tightening trigger embeddings, since in this case the watermark task is more independent from the normal task, making the optimization more stable. Overall, the ablation study has demonstrated the necessity of both the loss terms in our watermarking framework.
@@ V Conclusion
In this paper, we address the problem of IP protection of graph self-supervised encoders by proposing a primary watermarking scheme. By forcing the encoder to produce unique and compact embeddings for trigger instances, any downstream models integrated with the encoder always give similar prediction results for the samples in the trigger set, which constitutes an effective black-box watermarking strategy. Experiments have shown that the embedded watermark in the encoder can be reliably detected in the combined models in downstream tasks. The watermarking approach also meets other basic requirements for typical DNN watermarking. As the techniques of self-supervised learning and graph foundation models fast grow in graph domain, the IP protection of GSSL emerges as an increasingly important problem. In the future, we will endeavor to propose more advanced watermarking schemes for self-supervised models, such as resisting model extraction attack.

### v2

@@ IV-A3 Watermark Embedding
During encoder pre-training, we pre-train all the encoders using the aforementioned total loss to embed. The factors in the loss function $\lambda_{1}$, $\lambda_{2}$ are set as $5.0$, $1.0$ respectively by default. The trigger node set consists of 50 ego-graphs (centered at 50 randomly sampled nodes) by default. All the GSSL models are pre-trained using Adam optimizer [20] with an initialized learning rate $0.01$ for 2000 epochs.
IV-A4 Downstream Tasks
@@
Node classification. For node classification, we simulate the downstream classifier as a 2-layer multi-layer perceptron (MLP) with 256 neurons in each hidden layer. When training, we simulate the adversary by using 20 labeled nodes per class to train the downstream classifier. In addition to that, 500 nodes are used for validation and 1000 nodes are used for testing. The settings closely follow [3]. As commonly used in classification tasks, we use accuracy (ACC) to record the performance of the node classification task.
@@
Community detection. Following the literature of graph representation learning regarding the task of community detection, we first adopt PCA to reduce the dimensionality of node representations. Then, k-means clustering is used to cluster the node communities based on the representations. Here, we use normalized mutual information (NMI) for criterion.
TABLE I: The fidelity evaluation (clean model performance |||| watermarked model performance) of the watermarking method.
@@
Figure 2: The watermark performance w.r.t. model pruning with different pruning rates. Here, we record the curves of concentration scores of trigger embeddings (CS) and model normal performance in three downstream tasks (ACC, AUC and NMI). (a) GGD, Cora, (b) GGD, Citeseer.
@@ IV-B1 Fidelity
Fidelity means that the impact of performance degradation induced by watermarking should be low. Specifically, for tasks of node classification, edge prediction and community detection, fidelity is reflected by the difference of ACC, AUC and NMI with and without watermarking, respectively. The experimental results of fidelity evaluation are listed in Table I, where the normal performance of watermarked and non-watermarked models in downstream tasks are listed. It can be seen that, in most cases, the performance degradation caused by watermarking is negligible ($<1\%$), which is reasonable since the utility loss and the watermarking loss are optimized together. However, in some rare cases, the performance degradation is comparatively obvious, we argue that the unified factor settings in our experiments may not always lead to optimal solutions for all models and datasets. But overall, the training of the encoder with both utility loss and watermark loss can embed the watermark while maintaining the model performance. In practice, the defender can carefully fine-tune the factors.
TABLE II: The concentration score ($CS$) of the trigger predictions produced by watermarked models and non-watermarked models ($CS_{\text{wm}}\%|CS_{\text{clean}}\%)$.
@@
Figure 3: t-SNE visualization of the node embeddings generated by watermarked encoders. Scattered nodes in different colors represent the embeddings of nodes in different categories. Particularly, trigger embeddings are in black. (a) GGD, cora, (b) GGD, citeseer, (c) GraphCL, cora, (d) GraphCL, citeseer.
Figure 4: The watermark performance (concentration scores) w.r.t. different $\lambda_{1}$ and $\lambda_{2}$ settings. (a) GGD, Cora, (b) GGD, Citeseer.
@@ IV-B2 Transferability
Transferability means that watermarked graph encoder can be verified with models integrated with the encoder. In the proposed method, it means that the concentration score of the watermarked model is larger than a predetermined threshold $\tau$. In our experiments, we set the $\tau$ to be 0.7, which can well discriminate between watermark and clean models. The transferability evaluations are recorded in Table II, from which we can see that the concentration scores ($CS_{\text{WM}}$) are generally higher than 0.8 for watermarked models, demonstrating the effectiveness of the watermarking method.
@@ IV-B3 Uniqueness
Uniqueness means that the watermark can be extracted from only watermarked models. To show the uniqueness of the watermark, we list the concentration scores of trigger set representations produced by clean models in Table II. From the table we can observe that, the concentration scores of prediction results produced by watermarked models are higher than that of non-watermarked clean models by a large margin. For watermarked models, the predictions of trigger set generally exhibit high concentration score greater than the threshold $\tau=0.7$. In contrast, for non-watermarked models, since such models have not learned to produce similar embeddings for trigger samples, the result distribution is more uniform, which yields low concentration scores.
@@ IV-B4 Robustness
As a primary approach, we mainly focus on testing the robustness against model parameter pruning since it is the most common way to modify neural networks. Specifically, we record the change of concentration score and normal performance after pruning $\alpha\%$ of parameters with largest magnitude. As seen in Figure 2, generally the concentration score of watermarked model degrades simultaneously with the model normal performance of different tasks as the pruning rate increases. The phenomenon indicates that it is hard for the adversary to remove the watermark without degrading the model normal performance.
@@ IV-B5 Visualization
To further explain the effectiveness of the proposed method, we show the t-SNE visualization of the embeddings produced by watermarked encoder in Figure 3. It can be seen that, the normal nodes are well represented by embeddings while the trigger embeddings form a compact cluster (black dots in the figure) that is comparatively distant from normal samples. The t-SNE visualization clearly shows that the design of watermark loss successfully accomplishes our main objective and thus results in effectiveness watermarking for GSSL encoders.
@@ IV-C Ablation Study
To validate the necessity of the modules in the watermarking scheme, we perform an ablation study. Since the watermark is embedded by the watermark loss, so we mainly focus on studying the necessity of loss terms, including $L_{\text{in}}$ and $L_{\text{ext}}$, which are scaled by $\lambda_{1}$ and $\lambda_{2}$, respectively. Here, we focus on node classification and calculate the concentration scores of classification results of models with encoders pre-trained with different loss configurations using different $\lambda_{1}$ and $\lambda_{2}$. The results are shown in Figure 4. It can be seen that, generally speaking, $\lambda_{1}$ plays a main role with respect to the resulting concentration score, which is reasonable since $L_{\text{ext}}$ clusters the embeddings of trigger nodes. But we can also observe the effect of $L_{\text{ext}}$ scaled by $\lambda_{2}$, since in some cases, only using $L_{\text{int}}$ (by setting $\lambda_{1}$ to be zero) leads to sub-optimal results. Overall, through the ablation study, we find that $L_{\text{int}}$ and $L_{\text{ext}}$ both contribute to the optimization of the encoder during watermark embedding.
@@ V Conclusion
In this paper, we address the problem of IP protection of graph self-supervised encoders by proposing a primary watermarking scheme. By forcing the encoder to produce unique and compact embeddings for trigger instances, any downstream models integrated with the encoder always give similar prediction results for the samples in the trigger set, which constitutes an effective black-box watermarking strategy. Experiments have shown that the embedded watermark in the encoder can be reliably detected in the combined models in downstream tasks. The watermarking approach also meets other basic requirements for typical DNN watermarking. As the techniques of self-supervised learning and graph foundation models rapidly grow in the graph domain, the IP protection of GSSL emerges as an increasingly important problem. In the future, we will endeavor to propose more advanced watermarking schemes for self-supervised models, such as resisting model extraction attack.